OpenStudy (anonymous):

Find GCF and LCM 420, 504

OpenStudy (anonymous):

420 = 7*3*2*5*2 504=2*2*2*7*3*3 GCF = 7*2*3=84 LCM = (product of numbers)/hcf = (420*504)/84 = 2520

OpenStudy (across):

Factors of 420: 2^2, 5, 3, 7 Factors of 504: 2^3, 3^2, 7 GCF: (2^3)(3^2)(5)(7) = 2,520 LCM: (2^2)(3)(7) = 84